OpenStudy (anonymous):

determine whether the vectors <2, -1> and <-2, -5> are orthogonal

OpenStudy (anonymous):

if the dot product is 0 then they are orthogonal

OpenStudy (anonymous):

hint dot product of \[<x _{1},y _{1}>\text{ & }<x _{2},y _{2}>=x _{1} \times x _{2} + y _{1} \times y _{2} \]

OpenStudy (anonymous):

so the dot product for this would be (2*-2)-(1*-5) =-4+5=1 so not orthogonal

OpenStudy (anonymous):

2(-2) + (-1)(-5) = -4 + 5 = 1, excellent

OpenStudy (anonymous):

YAY thanks!!!

OpenStudy (anonymous):

you're welcome